# The Impact of Meteorological Factors on Crop Price Volatility in India: Case studies of Soybean and Brinjal

### v1

@@ 3.1 EGARCH Model for Price Volatility Estimation
We begin our analysis by fitting the Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH) model to the log-return of soybean and brinjal price data across Madhya Pradesh and Odisha respectively.
@@ 3.1.1 Methodology
The EGARCH model, an extension of the GARCH family, is particularly suited for financial time series with asymmetric volatility. It captures the so-called "leverage effect," where negative shocks to the market result in larger increases in volatility compared to positive shocks of the same magnitude.The EGARCH model is defined by the following equations for the conditional variance $\sigma_{t}^{2}$:
$\ln\sigma_{t}^{2}=\omega+\sum_{i=1}^{p}\alpha_{i}\left(\left|\frac{\epsilon_{t-i}}{\sigma_{t-i}}\right|-\sqrt{\frac{2}{\pi}}\right)+\sum_{j=1}^{o}\gamma_{j}\frac{\epsilon_{t-j}}{\sigma_{t-j}}+\sum_{k=1}^{q}\beta_{k}\ln\sigma_{t-k}^{2}$
@@
  • $\omega$: The constant term representing the baseline level of the log conditional variance.
  • $\alpha_{i}$ (for $i=1,\dots,p$): These parameters capture the impact of the magnitude of past shocks on current volatility. The term
@@ 3.1.2 Results
Figure 4: Conditional Volatility: The plots represent the conditional volatility of monthly log returns of Soybean and Brinjal prices from 2012 to 2024 respectively.
Figure 4 illustrates the conditional volatility of monthly log returns using the EGARCH model for Madhya Pradesh and Odisha, effectively capturing the volatility patterns in the prices of soybean and brinjal in these states. From the figure, we observe that soybean experienced a significant spike in volatility during late 2021 when the impact of COVID-19 pandemic was still fresh. As it is an export-oriented oilseed, soybean was heavily affected by the aftereffects of the lockdown. These factors impacted their supply-demand chain resulting in extreme volatility in their prices, which is evident from Fig. 2 as well. Significant volatility spike in soybean price is observed during 2013 as well which was due to weakness in rupee and the soybean import demand from Iran [25]. On the other hand, as brinjal is a vegetable which is mainly consumed in the domestic market, it does not show such drastic spike in volatility. This shows that the sensitivity towards various events are different for different crops. Table 2 contains the parameter estimates obtained from the EGARCH model fitted on crop price volatility.
Table 2: Parameter Estimates for Madhya Pradesh and Odisha
Parameter Madhya Pradesh Odisha
@@
$\mu$ (Mean) 0.00063 0.009570 0.010211 1.21e-10
$\omega$ (Constant) -0.663758 0.206315 -0.433305 1.90e-11
$\alpha_{1}$ (ARCH Coefficient) 0.625088 0.234577 -0.569581 2.57e-10
$\alpha_{2}$ (ARCH Coefficient) -0.012992 0.102649 - -
$\alpha_{3}$ (ARCH Coefficient) -0.862315 0.320954 - -
$\gamma_{1}$ (Leverage Effect) -0.004637 0.210729 0.012321 1.14e-11
$\gamma_{2}$ (Leverage Effect) -0.071931 0.222575 - -
$\gamma_{3}$ (Leverage Effect) 0.490231 0.173930 - -
$\beta_{1}$ (GARCH Coefficient) 0.091624 0.215075 0.398459 1.01e-12
$\beta_{2}$ (GARCH Coefficient) 0.801353 0.189453 0.038338 5.31e-12
$\beta_{3}$ (GARCH Coefficient) - - 0.430995 1.33e-11
3.2 Cross-Correlation and Granger-causality
@@ 3.2.1 Methodology
$\text{CCF}(k)=\frac{\sum_{t=k+1}^{T}(X_{t}-\bar{X})(Y_{t-k}-\bar{Y})}{\sqrt{\sum_{t=k+1}^{T}(X_{t}-\bar{X})^{2}\sum_{t=k+1}^{T}(Y_{t-k}-\bar{Y})^{2}}}$
@@
Granger-causality test is a statistical test to identify the cause and effect of one time series on another [26]. We estimate the Granger-causality between volatility and past meteorological variables to analyze the cause of the meteorological variables on the volatility. In order to estimate the cause of time series $Y$ on time series $X$, a regression is performed. $X$ is modeled from it’s own lagged values and the lagged values of $Y$. Further, we evaluate the significance of coefficient linked with $Y$ to check if they are useful for forecasting $X$
$X_{t}=\alpha_{0}+\alpha_{1}X_{t-1}+...+\alpha_{k}X_{t-k}+\beta_{1}Y_{t-1}+...+\beta_{k}Y_{t-k}+\epsilon_{t}$ (1)
where k is the no. of lags, $\alpha$ and $\beta$ are the coefficients of the lagged values of $X$ and $Y$, respectively, and $\epsilon$ is the prediction error.
Granger-causality employs hypothesis testing to evaluate the significance of the coefficients by computing the $p$ value [26]. The null hypothesis states that $X$ does not cause $Y$ indicating that the lagged values does not hold any information to improve the forecasting. If the $p$ value is greater than 0.05 we accept the null hypothesis. However, we reject the null hypothesis if the $p$ value is less than 0.05.
@@ 3.2.2 Results
The cross-correlation analysis measures how volatility is related to different lagged values of meteorological variables. Figure 5 shows the cross-correlation between volatility with precipitation and temperature, respectively. We have calculated the cross-correlation for different lags from -60 to +60. The vertical dashed line signifies the correlation at lag 0. From the correlation plot between volatility and precipitation, we observe a noticeable rise in correlation at lag 3 and 4. This suggests that there is a positive relationship between volatility and precipitation at lag 3 and 4. Similarly, a relatively high correlation is observed between the volatility and maximum temperature at lag 4. This shows that around lag 3/4 the correlation between volatility and meteorological variables are relatively high. This high correlation is repeating at a lag of 12 which shows the annual cycle as each lag signifies one month period. It would be interesting to know the cause-effect relationship of these variables. It is interesting to observe that there is no such seasonal correlation between volatility and meteorological variables in the case of brinjal in Odisha, as seen in fig. 6. From cross correlation, no conclusive relation is obtained for the case of brinjal.
@@
Refer to captionCross-CorrelationLag
Figure 5: Cross-Correlation Analysis. Cross-correlation analysis with 60 lags for Madhya Pradesh (MP): First is for cross-correlation between Volatility and Precipitation, and second is for cross-correlation between Volatility and Maximum Temperature with varying time lags.
After the estimation of CCF, it is important to analyze the cause-effect relation between volatility and the meteorological variables. In order to check for a causal relation between volatility and meteorological variables, we conducted a Granger-causality test for both price time series. We first checked the stationarity using Kwiatkowski-Phillips-Schmidt-Shin (KPSS) and Phillips-Perron tests for volatility, maximum temperature, and precipitation for both cases. We found that both time series are stationary. Hence, we estimated the Granger causality to test whether maximum temperature and precipitation have an effect on the volatility. We found that for Madhya Pradesh, there exists a causal relation between the volatility and meteorological variables at lag 3. The lag of 3 months is consistent with the sowing and harvest of the soybean crop. It is sown during the onset of the monsoon in the month of June/July and harvested during October. However, in the case of Odisha, we see no causal relation between the volatility and meteorological variables. This may be due to the fact that brinjal is grown all year round and is not specific to certain conditions. However, the meteorological variables play an important role in the yield and price of brinjal. So, we use the meteorological variables as inputs to forecast the volatility using ARIMAX and LSTM.
@@
Refer to captionCross-CorrelationLag
Figure 6: Cross-Correlation Analysis. Cross-correlation analysis with 60 lags for Odisha: the first plot is for cross-correlation between Volatility and Precipitation, and the second is for cross-correlation between Volatility and Maximum Temperature, each with varying time lags.
3.3 Predictive Modeling (SARIMAX and LSTM)
@@ 3.3.1 Methodology
The third stage of our analysis focuses on developing predictive models that forecast price volatility by incorporating meteorological inputs. To achieve this, we employ two distinct modeling approaches: the SARIMAX (Seasonal Autoregressive Integrated Moving Average with Exogenous variables) model and the LSTM (Long Short-Term Memory) neural network model. Each method offers unique advantages in capturing the temporal dynamics and complex relationships between climatic variables and agricultural price volatility.
SARIMAX Model
@@
$\Phi_{P}(L^{s})\phi_{p}(L)(1-L)^{d}(1-L^{s})^{D}\sigma_{t}=\Theta_{Q}(L^{s})\theta_{q}(L)\varepsilon_{t}+\mathbf{x}_{t}\boldsymbol{\beta}$
@@
The SARIMAX model captures the linear relationships between the dependent variable (price volatility) and both its own past values and external meteorological variables while also accounting for seasonal patterns. By incorporating these exogenous variables, the model can predict future volatility based on past patterns in the data and concurrent climate conditions. By incorporating these exogenous variables, the model can predict future volatility based on past patterns in the data and concurrent climate conditions. The model fitting involves selecting appropriate values for $p$, $q$, $d$, $P$, $Q$ and $D$ where $d$ and $D$ represent the differencing and seasonal differencing order to make the time series stationary when needed. To determine the optimal model parameters, we optimized the Akaike Information Criterion (AIC), which helped balance model complexity with predictive accuracy.
LSTM Model
To capture non-linear and complex dynamics between climatic variables and price volatility, LSTM model is used. The LSTM model is a type of recurrent neural network (RNN) that is designed to handle long-term dependencies in sequential data, making it well-suited for time series forecasting in agricultural markets, where the effects of weather patterns on prices may exhibit delayed interactions.
@@
The Forget Gate in an LSTM model decides what information should be discarded from the memory cell, computed as
@@
where $f_{t}$ is the forget gate vector, $\sigma$ is the sigmoid activation function, $W_{f}$ are the weight matrices, $h_{t-1}$ is the previous hidden state, $x_{t}$ is the current input, and $b_{f}$ is the bias. The Input Gate determines what new information should be added to the memory, calculated as
@@ 3.3.2 Results
This section contains the forecast results of the conditional volatility using the SARIMAX and LSTM models. We use the meteorological variables (i.e., precipitation and maximum temperature) as independent variables. Each model is tested on two crops from different states: soybean from Madhya Pradesh, and brinjal from Odisha. We use mean absolute percentage error (MAPE) to estimate the predictive accuracy of the models.
Figure 7 represents the conditional volatility plots for soybean and brinjal, respectively. The solid lines, brown for soybean and purple for brinjal, show the conditional volatility obtained from the EGARCH model. The dashed lines, orange for soybean and green for brinjal, are the volatility predicted from the SARIMAX model during the testing period. We obtained a MAPE of 0.55 and 0.19 for soybean and brinjal, respectively. The result shows that forecasting error is higher for the price volatility of soybean. This may be because the volatility during the testing period (Sep 2019 - Oct 2024) was more dominated by the external factors such as COVID-19 than the meteorological factors. As earlier discussed as soybean is an export-oriented crop it is impacted by global disruptions. On the other hand brinjal is not affected by the disruptions hence, we obtain a lower MAPE.
@@
Figure 7: Comparison of SARIMAX Forecasting Models. The top plot illustrates the conditional volatility of log returns for Soybean prices in Madhya Pradesh (MP), while the bottom plot shows the conditional volatility of log returns for Brinjal prices in Odisha. In the first plot, the orange line represents the SARIMAX model’s prediction for Soybean price conditional volatility during the testing period from September 2019 to October 2019. Similarly, in the second plot, the green line represents the SARIMAX prediction for Brinjal price conditional volatility over the same testing period.
Similarily, in figure 8 the solid lines, brown for soybean and purple for brinjal, show the conditional volatility obtained from the EGARCH model. The dashed lines, orange for soybean and green for brinjal, are the volatility predicted from the LSTM model during the testing period. We obtained a MAPE of 0.45 and 0.21 for soybean and brinjal, respectively. Due to similar reasons as stated above, the forecasting error may be higher for the price volatility of soybean.
@@
Figure 8: LSTM Forecasting Models Comparison.The top plot illustrates the conditional volatility of log returns for Soybean prices in Madhya Pradesh (MP), while the bottom plot shows the conditional volatility of log returns for Brinjal prices in Odisha. In the first plot, the orange line represents the LSTM model’s prediction for Soybean price conditional volatility during the testing period from December 2019 to October 2019. Similarly, in the second plot, the green line represents the LSTM prediction for Brinjal price conditional volatility over the same testing period.
@@
Madhya Pradesh Soybean SARIMAX 0.55
LSTM 0.45
Odisha Brinjal SARIMAX 0.19
LSTM 0.21
Table 3: RMSE and MAPE Comparison for Different Models
4 Discussions
We obtain various implications from the results obtained from our work. The analysis of EGARCH shows that the two crops exhibit significantly different volatility patterns due to the nature of the product, their market demand, and their sensitivity towards external factors. We found that soybean is more sensitive, as it is an export-oriented product that is susceptible to global disruptions. On the other hand, brinjal, being a vegetable, is mostly consumed in the domestic market, and hence, it is not volatile despite global disruptions. This is also supported by the statistical analysis, which showed that for soybean, the skewness and kurtosis indicate a higher chance of extreme movements than brinjal. From these results, we could say that policy for export-specific crops should be framed keeping into consideration their sensitivity towards external factors.
The CCF results indicated that the volatility of soybean is more correlated to the meteorological variables at lags 3 and 4, indicating that these variables impact the price volatility. The lag of 3 and 4 is also consistent with the sowing period (June/July) and harvesting period (October) of the soybean. This is further justified by the Granger-causality test, which shows a causal relationship between the meteorological variables and volatility at lag 3. This clearly shows that soybean is quite sensitive to weather conditions. Brinjal, on the other hand, shows no significant relation with the meteorological variables. Brinjal is grown throughout the year and does not have strict seasonal cycles; hence, the low correlation and absence of causal relation are meaningful. However, meteorological variables play a crucial role in the production and pricing of brinjal.
The application of models such as SARIMAX and LSTM will help to forecast future price volatility based on meteorological variables. Also, it will help us choose the better model for such analysis. The supremacy of one model is hard to find from our analysis as for soybean LSTM performs better, and for brinjal, LSTM performs better based on the MAPE. However, we clearly see that the overall MAPE for brinjal is lower than for soybean. This is consistent with the results obtained from volatility, which indicated that soybean is more volatile around 2020 due to the COVID-19 effect. This clearly shows that apart from the meteorological variables, which are important for forecasting, other factors also affect the price volatility and should be included in the model to improve the accuracy.
This work will have more implications, particularly in areas that are more affected by climate change. The forecasts will help farmers optimize their sowing and harvesting time. They can make rational decisions by considering how the weather patterns will impact prices. They can also make a beneficial decision regarding crop rotation and allocations. In general terms, implementing meteorological factors to assess agricultural performance could help to understand and reduce price volatility that may result in economic stability.
5 Summary
Our study highlights the impact of meteorological variables on the price volatility of soybean and brinjal. Price volatility is estimated using the EGARCH model which shows that soybean in comparision to brinjal shows more sensitivity towards global disruptions like COVID-19. The export-oriented nature of soybean is the main reason for such sensitivity. On the other hand, as brinjal is a vegetable mainly focused on domestic market is less sensitive to external disruptions. This result also aligns with the statistics of log-return which shows that soybean has higher skewness and kurtosis indicating a higher change of extreme movements. Further, CCF and Granger-causality test shows that price volatility of soybean, which is a kharif crop, is significantly impacted by the meteorological variables at lag 3 and 4. This lag is consistent with the period between sowing and harvesting of soybean. We found that brinjal is not significantly impacted by the meteorological variables. The finding is also consistent as it is cultivated almost all year round and is not dependent on a particular season for cultivation. Lastly, SARIMAX and LSTM is applied to forecast the conditional volatility using the meteorological variables. The forecasting of price volatility of soybean is less accurate than brinjal. This may be because of the higher volatility in soybean during 2020.
Our findings could be beneficial in areas more vulnerable to climate change. It may also help policy-makers to make policies which are robust to external disruptions. Also, in future other variables in addition to meteorological variables may be incorporated in the model to understand the price dyanmices better and also improve the accuracy.

### v3

@@ 4.1 EGARCH Model for Price Volatility Estimation
We begin our analysis by fitting the Exponential Generalized Autoregressive Conditional Heteroskedasticity (EGARCH) model to the log-returns of soybean and brinjal price data across Madhya Pradesh and Odisha respectively.
@@ 4.1.1 Methodology
The EGARCH model, an extension of the GARCH family, is particularly suited to capture asymmetric volatility in time-series data. It models the conditional variance, $\sigma_{t}^{2}$, as an asymmetric function of lagged disturbances defined by:
$\ln\sigma_{t}^{2}=\omega+\sum_{i=1}^{p}\alpha_{i}\left(\left|\frac{\epsilon_{t-i}}{\sigma_{t-i}}\right|-\sqrt{\frac{2}{\pi}}\right)+\sum_{j=1}^{o}\gamma_{j}\frac{\epsilon_{t-j}}{\sigma_{t-j}}+\sum_{k=1}^{q}\beta_{k}\ln\sigma_{t-k}^{2}.$
@@
  • $\omega$: The constant term represents the baseline level of the log conditional variance.
  • $\alpha_{i}$ (for $i=1,\dots,p$): These parameters capture the impact of past shocks on current volatility. The term
@@ 4.1.2 Results
Figure 5: Conditional Volatility. The brown and purple lines represent the conditional volatility of monthly log returns for soybean and brinjal prices respectively.
Figure 5 illustrates the conditional volatility of monthly log returns using the EGARCH model for soybean and brinjal prices. A notable spike in soybean price volatility is observed in late 2021, coinciding with the lingering effects of COVID-19 pandemic. As an export-oriented oilseed, soybean was heavily affected by the aftereffects of the lockdown, impacting its supply chain resulting in extreme volatility in their prices. Being a kharif crop, it is typically sown during the onset of monsoon in the months of June-July and harvested during October. Hence, volatility spikes are observed primarily during specific periods, likely due to harvest and supply fluctuations. Outside these periods, price volatility movements remain relatively stable due to consistent supply. In contrast, brinjal does not have a specific harvesting season therefore exhibits persistent volatility throughout the year. A sharp increase and decline in brinjal prices can be observed during 2014 (April-May), when delayed rainfall led to sharp increase in brinjal prices, followed by a rapid decrease as rainfall improved and supply stabilized[39].
Table 2 presents the parameter estimates obtained from the EGARCH model applied to crop price log returns.
Table 2: Parameter Estimates of EGARCH model for Soybean and Brinjal
Parameter Soybean Brinjal
@@
$\mu$ (Mean) 0.00877 0.00618 0.03656 8.559e-07
$\omega$ (Constant) -0.41060 0.11445 -0.22599 2.202e-12
$\alpha_{1}$ (ARCH Coefficient) 0.10250 0.16198 0.64441 1.002e-10
$\alpha_{2}$ (ARCH Coefficient) 0.15469 0.21614 -1.06278 3.357e-08
$\alpha_{3}$ (ARCH Coefficient) -0.37506 0.16464 0.15620 1.830e-09
$\gamma_{1}$ (Leverage Effect) 0.01845 0.13879 -0.06083 8.479e-07
$\gamma_{2}$ (Leverage Effect) 0.52901 0.17544 - -
$\gamma_{3}$ (Leverage Effect) -0.24236 0.17653 - -
$\beta_{1}$ (GARCH Coefficient) 0.92453 0.02249 0.29258 1.026e-10
$\beta_{2}$ (GARCH Coefficient) - - 0.62815 6.449e-10
4.2 Cross-Correlation and Causality
@@ 4.2.1 Methodology
$\text{CCF}(k)=\frac{\sum_{t=k+1}^{T}(X_{t}-\bar{X})(Y_{t-k}-\bar{Y})}{\sqrt{\sum_{t=k+1}^{T}(X_{t}-\bar{X})^{2}\sum_{t=k+1}^{T}(Y_{t-k}-\bar{Y})^{2}}},$
@@
Granger-causality test is a statistical test to identify the cause and effect of one time series on another [40, 41]. We estimate the Granger-causality between volatility and past meteorological variables to analyze the cause of the meteorological variables on the volatility. In order to estimate the cause of time series $Y$ on time series $X$, a regression is performed. $X$ is modeled from it’s own lagged values and the lagged values of $Y$. Further, we evaluate the significance of coefficient linked with $Y$ to check if they are useful for forecasting $X$
$X_{t}=\alpha_{0}+\alpha_{1}X_{t-1}+...+\alpha_{k}X_{t-k}+\beta_{1}Y_{t-1}+...+\beta_{k}Y_{t-k}+\epsilon_{t},$ (1)
where $k$ is the no. of lags, $\alpha$ and $\beta$ are the coefficients of the lagged values of $X$ and $Y$, respectively, and $\epsilon$ is the prediction error. Granger-causality employs hypothesis testing to evaluate the significance of the coefficients by computing the $p$ value [40, 41]. The null hypothesis states that $X$ does not cause $Y$ indicating that the lagged values does not hold any information to improve the forecasting.We reject the null hypothesis if the $p$ value is less than 0.05.
@@ 4.2.2 Results
Refer to captionCross-CorrelationMonthly Lag
Figure 6: Cross-Correlation Analysis of soybean price volatility in Madhya Pradesh. The plot represents the cross-correlation analysis with 24 monthly lags for soybean price volatility; the first plot represents the cross-correlation between volatility and precipitation, while the second plot represents the cross-correlation between volatility and maximum temperature.
Figure 6 and  7 showcase the cross-correlation of crop price volatility and meteorological variables for soybean in Madhya Pradesh and brinjal in Odisha respectively. In both cases, we observe a cyclical relationship highlighting the presence of seasonality. The periodic nature of the correlation suggests that weather patterns influence price volatility in a recurring manner. However, the strength of the correlation remains relatively low, with values not exceeding 0.25 or dropping below -0.25 for soybean and remaining within the range of 0.1 to -0.1 for brinjal. This indicates that while meteorological variables exhibit a seasonal impact on volatility, their direct influence is not strongly pronounced, particularly in the case of brinjal.
@@
Refer to captionCross-CorrelationMonthly Lag
Figure 7: Cross-Correlation Analysis of brinjal price volatility in Odisha. The plot represents the cross-correlation analysis with 24 monthly lags for brinjal price volatility; the first plot represents the cross-correlation between volatility and precipitation, while the second plot represents the cross-correlation between volatility and maximum temperature.
After estimating the cross-correlation function, it is important to analyze the cause-effect relationship between volatility and meteorological factors. In order to check for a causal relation between volatility and meteorological variables, we conducted a Granger-causality test for both price volatility time series. The stationarity of volatility, maximum temperature, and precipitation was examined using the Kwiatkowski-Phillips-Schmidt-Shin (KPSS) test for both cases. The results confirmed that the time series for soybean price volatility is stationary, whereas brinjal price volatility is not. To address this, the first difference of brinjal price volatility was taken to achieve stationarity, after which we proceeded with the Granger causality test to examine whether maximum temperature and precipitation have an effect on price volatility.
The analysis did not reveal any significant casual relationship between meteorological variables and soybean price volatility in Madhya Pradesh. This can be attributed to the fact that soybean is a globally traded commodity and its prices are largely influenced by international demand and supply conditions. As a result local climate change factors in Madhya Pradesh have minimal impact, which is reflected in the absence of a Granger causal relationship between soybean price volatility and meteorological factors. In contrast, brinjal, a vegetable primarily produced in Odisha, shows a different pattern. In this case, price volatility is related to local demand and supply, which are more sensitive to regional weather conditions. We observe a Granger casual relationship between precipitation and differenced price volatility at lag 6 and 7. This indicates that rainfall patterns significantly influence the market stability of brinjal in Odisha.
4.3 Predictive Modelling
@@ 4.3.1 Methodology
The next stage of our analysis focuses on developing predictive models that forecast price volatility by incorporating meteorological inputs. To achieve this, we employ two distinct machine learning approaches[42, 43]: the SARIMAX (Seasonal Autoregressive Integrated Moving Average with Exogenous variables) model and the LSTM (Long Short-Term Memory) neural network. Each method offers unique advantages in capturing the temporal dynamics and complex relationships between climatic variables and agricultural price volatility.
SARIMAX Model:
@@
$\Phi_{P}(L^{s})\phi_{p}(L)(1-L)^{d}(1-L^{s})^{D}\sigma_{t}=\Theta_{Q}(L^{s})\theta_{q}(L)\varepsilon_{t}+\mathbf{x}_{t}\boldsymbol{\beta},$
@@
The model captures the linear relationships between the dependent variable (price volatility) and both its own past values and external meteorological variables while also accounting for seasonal patterns. By incorporating these exogenous variables, the model can predict future volatility based on past patterns in the data and concurrent climate conditions. The model fitting involves selecting appropriate values for $p$, $q$, $d$, $P$, $Q$ and $D$, where $d$ and $D$ represent the differencing and seasonal differencing order respectively, to make the time series stationary when needed. To determine the optimal model parameters, we optimized the Akaike Information Criterion (AIC)[43], which helped balance model complexity with predictive accuracy.
LSTM Model:
To capture non-linear and complex dynamics between climatic variables and price volatility, an LSTM model is used. The LSTM model is a type of recurrent neural network (RNN)[42, 43] that is designed to handle long-term dependencies in sequential data, making it well-suited for time series forecasting in agricultural markets, where the effects of weather patterns on prices may exhibit delayed interactions.
@@
The Forget Gate in an LSTM model regulates which past information is retained or discarded in the memory cell, computed as
@@
where $f_{t}$ is the forget gate vector, $\sigma$ is the sigmoid activation function, $W_{f}$ are the weight matrices, $h_{t-1}$ is the previous hidden state, $x_{t}$ is the current input, and $b_{f}$ is the bias. The Input Gate determines what new information should be updated into the memory, calculated as
@@ 4.3.2 Results
Figure 8 compares the conditional volatility obtained from the EGARCH model with the SARIMAX-predicted volatility for soybean and brinjal, respectively. The results indicate that the forecasting error is higher for the price volatility of soybean. This may be because the volatility during the testing period (Sep 2019 - Oct 2024) was more driven by the external shocks such as COVID-19 than meteorological factors. As previously discussed, soybean, being an export-oriented crop, is more susceptible to global disruptions. In contrast, brinjal, which is primarily domestically consumed, is less affected by such external influences.
@@
Figure 8: Comparison of SARIMAX Forecasts. The brown and purple lines represent the conditional volatility of log-returns for soybean prices in Madhya Pradesh and brinjal prices in Odisha respectively. In first plot, the orange dashed line represents the SARIMAX prediction for soybean price volatility during the testing period from May 2020 to October 2024. Similarly, in second plot, the green dashed line represents the SARIMAX prediction for brinjal price volatility over the testing period of September 2019 to October 2024.
Similarly, Fig. 9 compares the conditional volatility obtained from the EGARCH model with the LSTM-predicted volatility for soybean and brinjal, respectively. As discussed earlier, the higher forecast error for soybean price volatility is likely due to external disruptions such as COVID-19 playing a more dominant role, whereas brinjal, being less affected by global shocks, exhibits a lower prediction error.
@@
Figure 9: Comparison of LSTM Forecasts.The brown and purple lines represent the conditional volatility of log returns for soybean prices in Madhya Pradesh and brinjal prices in Odisha respectively. In first plot, the orange dashed line represents the LSTM model’s prediction for soybean price volatility during the testing period from May 2020 to October 2024. Similarly, in second plot, the green dashed line represents the LSTM prediction for brinjal price volatility over the testing period of September 2019 to October 2024.
@@
Madhya Pradesh Soybean SARIMAX 0.48
LSTM 0.33
Odisha Brinjal SARIMAX 0.26
LSTM 0.24
Table 3: MAPE Comparison for SARIMAX and LSTM Models
4.4 Spatio-Temporal Volatility Estimation
4.4.1 Methodology
For the final step of our analysis, we study the spatial distribution of conditional volatility in soybean prices across Madhya Pradesh and brinjal prices across Odisha. As such, we construct a series of monthly volatility surfaces using a combination of spatial statistical methods.
For each month, we extract district-level crop price conditional volatility estimates and compute a spatial neighborhood structure using the k-Nearest Neighbors (KNN) algorithm [43], selecting the two nearest neighbors for each observation. This structure is converted to a symmetric binary adjacency matrix, representing first-order spatial contiguity–direct connections between geometrically proximate observations. This adjacency matrix serves as the basis for modelling spatial dependency using the Leroux Conditional Autoregressive (CAR) model [44], a Bayesian hierarchical framework that produces smoothed volatility estimates that balance local variation with geographic dependence. These smoothed values were then interpolated across a 0.01-degree resolution grid using Inverse Distance Weighting (IDW) [45]. This deterministic method assigns greater weight to closer observations, enabling the construction of continuous volatility surfaces that capture local heterogeneity while preserving the spatial dependence structure inferred from the CAR model.
4.4.2 Results
Figure 10 compares the district-wise conditional volatility of the monthly log returns estimated through the EGARCH model for soybean and brinjal prices respectively. It is apparent from the plots that certain districts exhibit substantially higher levels of price volatility compared to others. This suggests the presence of localised factors that can amplify price fluctuations in specific districts.
Refer to captionConditional Volatility
Refer to captionConditional VolatilityYear
Figure 10: District-wise Conditional Volatility. The brown and purple gradient lines represent the conditional volatility of soybean price in different districts of Madhya Pradesh and brinjal price in different districts of Odisha respectively. The black lines represent the state-wise conditional volatility.
Figure 11 presents the volatility surface for soybean price in Madhya Pradesh on October 2021 and brinjal price in Odisha on June 2020. This spatial analysis shows the interpolated distribution of volatility across the state, with higher peaks indicating places of greater uncertainty. The analysis was performed for each time step and the resulting frames were stitched into a GIF, which is displayed in the Crop Price Analysis app [46].
Figure 11: Volatility Surface Estimation using CAR model. The top row shows the 2D spatial distribution (left) and the corresponding 3D interpolated surface (right) of soybean price volatility in Madhya Pradesh for October 2021. The bottom row presents analogous plots for brinjal price volatility in Odisha for June 2020. The frames were taken from spatial plots available in the Crop Price Analysis app [46].
As mentioned before, soybean price volatility has been increasing across all districts in Madhya Pradesh post-2021 due to shocks caused by the Covid-19 pandemic. In the case of brinjal prices in Odisha, volatility has been noticeably greater post-2020 primarily due to erratic rainfall patterns affecting different districts unevenly [47]. These climate uncertainties, coupled with localised factors such as inadequate irrigation infrastructure, have further amplified fluctuations in production and market prices, leading to increased volatility at the district level.
5 Discussions & Way Forward
This study examines the impact of meteorological variables on the price volatility of soybean and brinjal, highlighting their differing sensitivities to external disruptions. The EGARCH model revealed that soybean exhibited higher volatility compared to brinjal, primarily due to its export-oriented nature, which made it more susceptible to global shocks such as the COVID-19 pandemic. In contrast, brinjal, being a domestically consumed vegetable, showed relatively stable price volatility patterns. Statistical analysis further supported this, as soybean displayed higher skewness and kurtosis, indicating a greater likelihood of extreme price movements.
To forecast price volatility, SARIMAX and LSTM models were applied using meteorological variables. While LSTM provided better accuracy for soybean, both models performed similarly for brinjal. However, the forecasting accuracy for soybean was lower overall, likely due to heightened volatility in 2020 caused by external disruptions beyond meteorological factors. This suggested that additional variables, such as global trade dynamics and policy interventions, could be incorporated to improve predictive accuracy. To further understand spatial variation in price volatility, the CAR model followed by IDW interpolation was applied monthly, to construct volatility surfaces for each state. These surfaces revealed clear differences in crop price volatility between districts, driven by both common shocks such as the COVID-19 pandemic and localized factors such as inadequate irrigation infrastructure and uneven market access.
Our findings have important implications, particularly for regions which are highly vulnerable to climate change. Improved volatility forecasts could help farmers optimize sowing and harvesting schedules, make informed crop allocation decisions, and mitigate economic risks associated with price fluctuations. An extension of this study can be found in our paper Predicting and Mitigating Agricultural Price Volatility Using Climate Scenarios and Risk Models[48], where we examine the complex interactions between crop price volatility, Minimum Support Prices(MSP), and insurance premiums. As agricultural markets become increasingly sensitive to both global shocks and local climate variability, a holistic approach becomes essential to safeguard farmer incomes and ensure food security. Future research could build on this work by integrating additional factors beyond meteorological variables to refine volatility predictions and better understand agricultural price dynamics. Furthermore, the methodology in this study can be replicated for other Indian states and translated into an interactive dashboard to support decision-making by policymakers, researchers, and other stakeholders.